OpenStudy (jenniferjuice):

change all of the following equations into slope- intercept form. 2x - 5y ≥ -10 x - y ≥ 3 2x - y < -4 x + 2y ≥ -8

OpenStudy (texaschic101):

slope intercept form is y = mx + b 1. 2x - 5y >= -10 subtract 2x from both sides -5y >= -2x - 10 Now divide both sides by -5 but remember in an inequality, when dividing by a negative, change the inequality sign y <= -2/5x - 10/-5 y <= -2/5x + 2 do you understand how I did this one ?

OpenStudy (texaschic101):

oops....I made a mistake...let me fix it -5y >= -2x - 10 divide by -5 and change inequality sign y <= -2x/-5 - 10/-5 y <= 2/5x + 2

OpenStudy (anonymous):

Solve for y Do you have questions on solving the inequalities?